For what Values of p will the expression p+8x-3x^2 always be positive?
Answer by venugopalramana(3286)

For what Values of p will the expression p+8x-3x^2 always be positive?
FOR NO VALUE OF P WILL IT BE ALWAYS POSITIVE.
QUADRATIC WILL ALWAYS HAVE ONE SIGN POSITIVE OR NEGATIVE IF ITS
DISCRIMINANT IS NEGATIVE.
THEN IT WILL ALWAYS BE +VE IF COEFFICIENT OF X^2 IS POSITIVE
OR IT WILL ALWAYS BE - VE. IF COEFFICIENT OF X^2 IS NEGATIVE,AS
IS THE CASE HERE.HENCE IF
D=8^2-4P(-3)<0...THAT IS 64-12P<0...THAT IS....P>64/12=16/3,THE FUNCTION WILL BE ALWAYS NEGATIVE.